# Quantum algorithms for calculating determinant and inverse of matrix and solving linear algebraic systems

### v3

@@ V Conclusions
The algorithm for computing the determinant of a square matrix is discussed in details. It uses multiqubit control-SWAP and control-Z operations. The depth of this algorithm is $O(N\log^{2}N)$. This algorithm utilized the subroutine realizing the completely antisymmetric tensor $\varepsilon$ (operator $W^{(1)}_{SA}$) that may have relevance in other algorithms. The outcome of this algorithm is split into two components: the absolute value of the determinant of the matrix $A$, denoted as $|\det(A)|$, and the argument of the determinant of the matrix $A$, denoted as $\arg(\det(A))$. The first part is obtained from the measurement of an ancilla, (although the probability of the needed outcome in such measurement decreases exponentially with $N$), while the argument of the determinant remains embedded in the probability amplitude of the output state $|\Psi^{(D)}_{out}\rangle$. The size of the subsystem $S$ encoding the matrix $A$ is $N\log N$, the size of ancillae $A_{k}$, $k=0,\dots,N-2$, is $\tilde{N}=O(N\log N)$. Both these sizes serve to estimate the size of the whole algorithm which is $O(N\log N)$.
It is remarkable that the $(N-1)\times(N-1)$-matrix inversion algorithm is essentially algorithm for calculating the determinant of the $N\times N$ matrix obtained by replacing the first row by the row of $1/\sqrt{N}$ and the first column by the column of $q$’s (up to the diagonal element). This special encoding provides the superposition of all products of matrix elements which are necessary for calculating minors included into the definition of the inverse matrix. The algorithm distributes these products with proper signs among the appropriate inverse-matrix elements which are encoded into the quantum state of certain subsystem thus forming the output state. The depth of the algorithm is $O(N^{2}\log N)$.
As an application of inversion we develop the algorithm for solving systems of linear algebraic equations. This algorithm consists of two blocks: matrix inversion and matrix multiplication. At that, the matrix inversion is most resource-consuming block which finally defines the depth of the algorithm (which is $O(N^{2}\log N)$), and probability of required ancilla state after measurement, $\sim 2^{-O(N\log N)}$. We have to notice that the depths of the algorithms for calculating the determinant and matrix inverse were obtained under assumption that the $U$-, $V$-operators in (13) can be applied in parallel. Otherwise we have to change the depths $O(N\log^{2}N)$ and $O(N^{2}\log\,N)$ to $O(N^{2}\log^{2}N)$ in both cases.
@@
One notable aspect is the use of ancillae to gather all the garbage generated during the calculation, which can then be discarded with a measurement of a one-qubit ancilla. The weak point of both algorithms is probability of obtaining the required state of ancilla after measuring. This probability is $\sim 2^{-O(N\log N)}$ in all three algorithms.
Combining the algorithm of matrix inversion with the algorithm of matrix multiplication we demonstrate the usage of two types of matrix encoding into the state of a quantum system. First one is the row-wise encoding used in the input of $N\times N$ matrix inversion algorithm (for simplicity, we assume that $N=2^{n}$). It requires $N$ subsystems of $\log N$ qubits each, and the state of $i$th subsystem encodes the elements of the $i$th row of the matrix. The result of matrix inversion is stored in the two subsystems of $\log N$ qubits each. These subsystems enumerate rows and columns respectively. This form of encoding of the inverse matrix allows to perform its multiplication with the column $\bm{b}$ properly encoded in the state of $\log N$-qubit subsystem to obtain the solution $\bm{x}$ of the linear system. This solution is a column whose elements are encoded into the state of a $\log N$-qubit subsystem $R$.
@@ VI.1 Matrix multiplication
We present an algorithm for multiplying an $N\times K$ matrix $A^{(1)}$ by a $K\times M$ matrix $A^{(2)}$ with the elements $A^{(i)}=\{a^{(i)}_{jk}\}$, $i=1,2$, assuming $N=2^{n}$, $K=2^{k}$, $M=2^{m}$ with positive integers $n$, $k$, $m$. The algorithm below is an alternative to that proposed in ZQKW_arxiv2023 . Both the depth and size of the algorithm are less than those parameters in ZQKW_arxiv2023 , although the estimations for them remains the same: depth is $O(\log K)$ and size is $O(\log K+\log N+\log M)$.
@@
where $I_{R_{2},j}$ is the identity operators acting on the $j$th qubit of the subsystem $R_{2}$. We have chosen the qubits of $C_{1}$ as control ones. However, if $C_{1}$ and $R_{2}$ are exchanged here and below in the description of the algorithm, it remains correct. All operators $W_{j}$, $j=1,\dots,k$ commute with each other. Applying the operator $W^{(1)}_{C_{1}R_{2}}$,
@@
to $|\Phi_{0}^{(m)}\rangle$ we obtain:
@@
selecting terms with $|0\rangle_{R_{2}}$ from the superposition state $|\Phi_{1}^{(m)}\rangle$ (the other terms are by definition garbage; their sum is denoted by $|g_{1}\rangle$ above). Now we apply the Hadamard transformations
@@
Figure 6: (a) Circuit for the matrix multiplication algorithm; (b) Notation for multiqubit CNOT. We omit superscript $(m)$ at $|\Phi_{k}\rangle$, $k=0,\dots,4$, and at $|\Psi_{out}\rangle$ for brevity.

### v4

@@ V Conclusions
The algorithm for computing the determinant of a square matrix is discussed in details. It uses multiqubit control-SWAP and control-Z operations. The depth of this algorithm is $O(N\log^{2}N)$. This algorithm utilized the subroutine realizing the completely antisymmetric tensor $\varepsilon$ (operator $W^{(1)}_{SA}$) that may have relevance in other algorithms. The outcome of this algorithm is split into two components: the absolute value of the determinant of the matrix $A$, denoted as $|\det(A)|$, and the argument of the determinant of the matrix $A$, denoted as $\arg(\det(A))$. The first part is obtained from the measurement of an ancilla, (although the probability of the needed outcome in such measurement decreases exponentially with $N$), while the argument of the determinant remains embedded in the probability amplitude of the output state $|\Psi^{(D)}_{out}\rangle$. The size of the subsystem $S$ encoding the matrix $A$ is $N\log N$, the size of ancillae $A_{k}$, $k=0,\dots,N-2$, is $\tilde{N}=O(N\log N)$. Both these sizes serve to estimate the space of the whole algorithm which is $O(N\log N)$ qubits.
It is remarkable that the $(N-1)\times(N-1)$-matrix inversion algorithm is essentially the algorithm for calculating the determinant of the $N\times N$ matrix obtained by replacing the first row by the row of $1/\sqrt{N}$ and the first column by the column of $q$’s (up to the diagonal element). This special encoding provides the superposition of all products of matrix elements which are necessary for calculating minors included into the definition of the inverse matrix. The algorithm distributes these products with proper signs among the appropriate inverse-matrix elements which are encoded into the quantum state of certain subsystem thus forming the output state. The depth of the circuit is $O(N^{2}\log N)$.
As an application of inversion we develop the algorithm for solving systems of linear algebraic equations. This algorithm consists of two blocks: matrix inversion and matrix multiplication. At that, the matrix inversion is most resource-consuming block which finally determines the depth of the algorithm (which is $O(N^{2}\log N)$), and probability of required ancilla state after measurement, $\sim 2^{-O(N\log N)}$. We have to note that the depths of the algorithms for calculating the determinant and matrix inverse were obtained under assumption that the $U$-, $V$-operators in (13) can be applied in parallel. Otherwise we have to change the depths $O(N\log^{2}N)$ and $O(N^{2}\log\,N)$ to $O(N^{2}\log^{2}N)$ in both cases.
@@
One notable aspect is the use of ancillae to gather all the garbage generated during the calculation, which can then be discarded with a measurement of a one-qubit ancilla. The weak point of both algorithms is the probability of obtaining the required state of ancilla after measuring. This probability is $\sim 2^{-O(N\log N)}$ in all three algorithms.
Combining the algorithm of matrix inversion with the algorithm of matrix multiplication we demonstrate the usage of two types of matrix encoding into the state of a quantum system. The first one is the row-wise encoding used in the input of algorithm for caculating the determinant of $N\times N$ matrix and matrix inversion algorithm (for simplicity, we assume that $N=2^{n}$). It requires $N$ subsystems of $\log N$ qubits each, and the state of $i$th subsystem encodes the elements of the $i$th row of the matrix. The result of matrix inversion is stored in the two subsystems of $\log N$ qubits each. These subsystems enumerate rows and columns respectively. This form of encoding of the inverse matrix allows to perform its multiplication with the column $\bm{b}$ properly encoded in the state of $\log N$-qubit subsystem to obtain the solution $\bm{x}$ of the linear system. This solution is a column whose elements are encoded into the state of a $\log N$-qubit subsystem $R$.
@@ VI.1 Matrix multiplication
We present an algorithm for multiplying an $N\times K$ matrix $A^{(1)}$ by a $K\times M$ matrix $A^{(2)}$ with the elements $A^{(i)}=\{a^{(i)}_{jk}\}$, $i=1,2$, assuming $N=2^{n}$, $K=2^{k}$, $M=2^{m}$ with positive integers $n$, $k$, $m$. The algorithm below is an alternative to that proposed in ZQKW_arxiv2023 . Both the depth and size of the algorithm are less than those parameters in ZQKW_arxiv2023 , although the estimations for them remains the same: the depth is $O(\log K)$ and space is $O(\log K+\log N+\log M)$ qubits.
Figure 5: (a) Circuit for the matrix multiplication algorithm. We omit superscript $(m)$ in $|\Phi_{k}\rangle$ $(k=0,\dots,4)$ and $|\Psi_{out}\rangle$ for brevity. (b) Notation for multiqubit CNOT.
@@
where $I_{R_{2},j}$ is the identity operators acting on the $j$th qubit of the subsystem $R_{2}$. We have chosen the qubits of $C_{1}$ as control ones. However, if $C_{1}$ and $R_{2}$ are exchanged here and below in the description of the algorithm, it remains correct. The operator $W^{(1)}_{C_{1}R_{2}}$,
@@
can be implemented by a depth-1 circuit, since all $W_{j}$ act on different qubits. Applying $W^{(1)}_{C_{1}R_{2}}$ to $|\Phi_{0}^{(m)}\rangle$ we obtain:
@@
thus selecting terms with $|0\rangle_{R_{2}}$ from the superposition state $|\Phi_{1}^{(m)}\rangle$ (the other terms are by definition garbage; their sum is denoted by $|g_{1}\rangle$ above). Now we apply the Hadamard transformations